CSePub Home Book Store Instructors Independent Authors Abo You have saved $23,695.00 for a three-year vacation to Germany. You will monthly compounding while you live in Munich. You will make your withdrawals at the beginning of the month keep your money invested in an account paying 804% APR with What is the balance on the account after the first month? (This is the balance prior to the withdrawal for the second month) Answer Format: Currency: Round to: 2 decimal places. Enter Answer Here Submit Answer +- Prev Problem t A ProblemsNext Problem Contact Us Privacy Policy Tems Copyright 2010 Conley Smth Publishing Adl Right reserved red 4:11 pm
Answer #1

Money in account = 23695
Number of Periods = 12*3 = 36
Rate per month = 8.04%/12
Per month withdrawal using annuity due formula = 23695/((1+8.04%/12)*(1+(1-8.04%/12)-36)/8.04%/12) = 738.01
Interest part of PMT = 23695*8.04%/12= 158.76

Balance at at end of month 1 = 23695 -(738.01-158.76) =23115.75

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well